# Probing Saturon-like Limits in QCD Systems

### v1

@@ I Introduction
The potential connection between the infrared behavior of the strong interaction and gravity has long been a subject of intense discussion. One of the most prominent manifestations of this idea is the application of the Anti-de Sitter/Conformal Field Theory (AdS/CFT) correspondence in strongly interacting theories [8, 9, 10, 11]. The rapid development of this framework has led to new interpretations in black hole thermodynamics, quantum field theory, and hadronic physics, particularly by relating certain black hole properties to entanglement and other features emerging in strong interaction studies [12, 13] (See others works related entropy [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 17, 28, 29, 30, 31, 32, 33].). As the fundamental theory of strong interactions, QCD provides a basis for exploring hadron structure and nuclear matter properties. Its infrared dynamics can also be investigated through the mapping of four-dimensional Minkowski field theories onto the boundary of a five-dimensional AdS spacetime. A notable recent example of a concrete connection between gravity and QCD is the so-called BCJ “double copy” relation [34], where gravitational perturbative amplitudes are constructed from QCD amplitudes supplemented by additional kinematic factors. As a distinctive subject of gravitational studies, black holes exhibit certain properties that may find correspondence in investigations of the strong interaction. As discussed in Ref. [12], it is possible to relate the behavior of systems near saturation to both highly occupied graviton states characteristic of black holes and the framework of the Color Glass Condensate (CGC) effective theory [35, 36].
@@
The central focus of this work is to explore saturation phenomena from the perspective of proton structure. It is well established that in high-energy hadrons, such as protons, the number of gluons increases rapidly as the Bjorken-xx–the longitudinal momentum fraction carried by gluons–becomes very small. However, when the gluon density reaches a sufficiently high level, nonlinear interactions among gluons, such as gluon-gluon recombination, become significant and inhibit the unbounded growth of gluon density. This phenomenon is referred to as gluon saturation. In other words, the gluon population cannot grow indefinitely without violating unitarity. A central question addressed in this work is whether a system constrained by unitarity, such as the saturated gluon system in a proton, can be connected to the concept of a saturon. This concept, introduced by Dvali et al., refers to highly occupied field configurations that reach maximal entropy while respecting unitarity. In Sec. II, we offer a concise review of QCD’s nonlinear small‑$x$ evolution equations and the phenomenon of gluon saturation within the proton. We also formalize physical quantities such as entropy and occupation number to serve as potential criteria for identifying a system as a saturon. In Sec. III, we present our computed observables and examine the relationship between high-gluon-occupancy protons and saturon behavior. Finally, we provide a summary and outlook. This study introduces a novel perspective on saturation in the proton system, suggesting that, near the saturation threshold, universal behavior can emerge independently of detailed microdynamics or initial conditions. Such a proton state represents a saturated condition comprised of highly occupied soft quanta, sufficient to form a cross section constrained by unitarity [3, 4, 12].
@@ II.1 Gluon saturation and small-$x$ evolution
The unbounded growth of high-energy gluons is not consistent with physical reality. In practice, as the number of gluons increasingly populates the phase space of the proton, spatial overlap among them becomes inevitable. When the overlapping gluons come sufficiently close, gluon recombination occurs, leading to saturation (See FIG. 1). This nonlinear effect is not accounted for within the framework of the BFKL equation. In fact, the earliest nonlinear corrections to the DGLAP equation were introduced in Refs. [43, 44] and are known as the GLR-MQ equation. Subsequently, the Balitsky-Kovchegov (BK) equation [45, 46, 47] was formulated based on the unitarity of the dipole scattering amplitude, incorporating nonlinear terms.
@@ II.2 Balitsky-Kovchegov equation
In this work, we employ the BK equation to characterize the saturation behavior of the proton and compare it with the criteria for a saturon system. The BK equation governs the evolution of the dipole amplitude with rapidity $Y=\ln 1/x$, and in momentum space, it can be expressed as [45, 46]
@@
The two forms of the equation are equivalent, implying that the dipole amplitude is a function of the evolution rapidity $Y(x)$ and the gluon transverse momentum $k$. Both formulations of the BK equation admit solutions in momentum space and can serve as inputs for calculating scattering amplitudes in diffractive processes as well as proton structure functions. Once the dipole amplitude is obtained from solving the BK equation, one can proceed to compute the unintegrated gluon distribution (UGD) and the gluon number density. Therefore, the solution to the BK equation captures the gluon evolution dynamics in the small-$x$ regime. For the nonlinear integro-differential form of Eq. (1), a Runge-Kutta scheme incorporating integral terms can be employed. In this work, we use the BKsolver algorithm as outlined in Ref. [48], which approximates the integral term using Chebyshev polynomials. This method proves highly effective and allows for both fixed and running coupling scenarios. For the differential form of the BK equation, it is also possible to obtain approximate analytical solutions. By transforming the equation into the form of the Fisher-Kolmogorov-Petrovsky-Piskunov (FKPP) equation [49, 50, 51, 52, 53, 54, 55], one can apply techniques developed for nonlinear partial differential equations to perform the analysis.
@@
For the case of running coupling (rc), the details of the treatment using the homogeneous balance method can similarly be found in Ref. [59]. In this work, we adopt the saturation scale prescription (SSP), whereby the running coupling scale is identified with the saturation scale $Q_{s}$. This choice reflects the characteristic momentum of gluons in the saturated regime and offers improved performance in describing the small-$x$ proton structure function $F_{2}$. The scale-dependent running coupling constant is expressed as [59]
@@
where $b=(11N_{c}-2N_{f})/12N_{c}$ with the number of color $N_{c}$ and number of flavor $N_{f}$, respectively.
@@
with the free parameter $c$. This is consistent with the reasoning behind Eq. (5), where, upon assuming a specific scale for the coupling constant, the analytical solution to the BK equation in the running coupling scenario can also be obtained using the homogeneous balance method. The explicit form is given as [59]
@@
A detailed discussion of the numerical solutions to the BK equation is beyond the scope of this work; here we briefly outline the method for extracting the saturation scale $Q_{s}^{2}$ from such solutions. Following the approach proposed in Refs. [48, 63], it is reasonable to define the transverse momentum corresponding to the saturated dipole amplitude at different $Y(x)$ as the saturation scale, i.e., by imposing the condition $\mathcal{N}(Y,Q_{s}^{2}(Y))=\mathcal{N}_{0}\sim\kappa$. Clearly, the value of $Q_{s}$ depends on the chosen threshold $\kappa$. Investigating how different values of $\kappa$ affect the extracted saturation scale is of interest, and the resulting behavior should be compared to the asymptotic analytical behavior reported in Refs. [50, 49]. Notably, for $\kappa=1$, the numerical solutions obtained using BKsolver [48] reproduce the expected asymptotic behavior well in both fixed and running coupling scenarios.
@@ II.3 Saturons, occupation number and entropy
$N_{g}(x,\bm{k}_{\perp},\bm{x}_{\perp})\equiv\frac{(2\pi)^{3}}{2(N_{c}^{2}-1)}\mathrm{\penalty 10000\ }x\frac{\mathrm{d}N_{\mathrm{gluon}}}{\mathrm{d}x\mathrm{\penalty 10000\ d}^{2}\bm{k}_{\perp}\mathrm{d}^{2}\bm{x}_{\perp}},$ (10)
where $\bm{k}_{\perp}$ and $\bm{x}_{\perp}$ are transverse momentum and impact parameter of gluon, respectively. It means the the UGD per unit transverse area. A frequently used simplified estimate takes the form:
$N_{g}(x,Q^{2})\simeq\frac{1}{Q^{2}}\frac{xG(x,Q^{2})}{\pi R^{2}},$ (11)
where $xG$ is the gluon distribution and can be written by the integration of UGD $\mathcal{F}$, $xG(x,Q^{2})=\int^{Q^{2}}\frac{dk^{2}}{k^{2}}\mathcal{F}(x,k^{2})$. The integral over the transverse impact parameter is generally interpreted as the transverse cross section of the boosted proton, $\pi R^{2}$, with proton radius $R=0.84$ fm. Regarding the explicit form of the UGD, one approach derived from [65] expresses it in terms of the dipole amplitude $\mathcal{N}$ obtained from the solution of the BK equation
$\mathcal{F}(x,k^{2})=\frac{N_{\mathrm{c}}R^{2}}{\alpha_{\mathrm{s}}\pi}\left(1-k^{2}\frac{\mathrm{d}}{\mathrm{d}k^{2}}\right)^{2}k^{2}\mathcal{N}(k^{2},x).$ (12)
@@
For simplicity, the lowest entropy state $S_{0}$ can be set to zero. All integration constants can be absorbed into quantities related to the gluon occupation number, which may be determined from experimental data or global QCD fits. It is important to emphasize that, in this analysis, the gluon occupation number (or gluon density) is obtained from the solution of the BK equation. Ultimately, the thermodynamic entropy of the system constructed via the Unruh effect and the emergent gluon mass takes the following form:
@@ II.4 From proton to nucleus: Glauber-Gribov-Mueller Scheme
where $Q_{s,p}=Q_{s}$ in Eq. (5) represents the nucleon saturation scale, $R_{A}$ is the nuclei radius with mass number $A$ and $r_{0}=1.2$ fm.
@@ III.1 Proton case
Through the mean-field approximation of the nonlinear evolution equation, namely the momentum-space BK equation, both its analytical and numerical solutions for the dipole amplitude can be related to the gluon distribution via the definition of the UGD. This connection allows the determination of the gluon occupation number and the corresponding thermodynamic entropy, as given in Eqs (11-17).
@@
The parameter choices for all results shown in the figure are specified in the main text and are taken from the corresponding references. The four subplots, arranged from left to right and top to bottom, represent respectively: the analytical solution of the BK equation with fixed coupling, the analytical solution with running coupling, the numerical solution with fixed coupling, and the numerical solution with running coupling. Each panel displays the resulting gluon occupation number and the associated entropy. It is observed that in all cases, even down to $x=10^{-7}$-well below the kinematic reach of current DIS experiments-neither the entropy nor the gluon occupation number approaches the inverse coupling constant, $1/\alpha_{s}$. One important reason for this is the presence of approximations in the construction of the entropy, and the omission of higher-order corrections in the evolution equations.
@@ III.2 Nuclear case
As seen in the figure, compared with the proton result, under the $x$-dependent growth of the entropy, the thermodynamic entropy of a highly occupied lead nucleus reaches the saturation criterion ($1/\alpha_{s}$) at around $x\sim 10^{-6}$. This indicates that the nucleus indeed enhances the entropy, mainly through the increase of the transverse radius after the high-energy boost (set by the nuclear radius) and the amplification of the nuclear saturation scale. However, this comparison has limitations, such as the choice of the dipole-amplitude form and uncertainties from nuclear geometric parameters. In this section we use fixed coupling and neglect detailed impact-parameter dependence and higher-order corrections, so the “nucleus reaches while the proton does not” criterion should be regarded as a conservative, trend-level conclusion.

### v2

@@ I Introduction
The potential connection between the infrared behavior of the strong interaction and gravity has long been a subject of intense discussion. One of the most prominent manifestations of this idea is the application of the Anti-de Sitter/Conformal Field Theory (AdS/CFT) correspondence in strongly interacting theories [8, 9, 10, 11]. The rapid development of this framework has led to new interpretations in black hole thermodynamics, quantum field theory, and hadronic physics, particularly by relating certain black hole properties to entanglement and other features emerging in strong interaction studies [12, 13] (See others works related entropy [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 17, 28, 29, 30, 31, 32, 33].). As the fundamental theory of strong interactions, QCD provides a basis for exploring hadron structure and nuclear matter properties. Its infrared dynamics can also be investigated through the mapping of four-dimensional Minkowski field theories onto the boundary of a five-dimensional AdS spacetime. Distinct from the holographic duality, another profound connection between gauge theories and gravity is found in the Bern-Carrasco-Johansson (BCJ) “double copy” construction [34]. While AdS/CFT relates a gauge theory to a gravity theory in a higher-dimensional space, the double copy framework establishes a direct relationship between scattering amplitudes in gauge theories and gravity within the same spacetime dimension, loosely summarized as “gravity = gauge $\times$ gauge”.
Regardless of the specific framework, black holes remain a focal point for understanding high-density gravitational states. Intriguingly, it has been proposed that the physics of black holes—specifically their saturation of entropy bounds—may share universal features with other high-occupancy systems. Motivated by such universality classes, recent works [12] have suggested a duality between the highly occupied graviton states in black holes and the saturated gluon states in the Color Glass Condensate (CGC) effective theory [35, 36].
@@
The central focus of this work is to explore saturation phenomena from the perspective of proton structure. It is well established that in high-energy hadrons, such as protons, the number of gluons increases rapidly as the Bjorken-xx–the longitudinal momentum fraction carried by gluons–becomes very small. However, when the gluon density reaches a sufficiently high level, nonlinear interactions among gluons, such as gluon-gluon recombination, become significant and inhibit the unbounded growth of gluon density. This phenomenon is referred to as gluon saturation. In other words, the gluon population cannot grow indefinitely without violating unitarity. A central question addressed in this work is whether a system constrained by unitarity, such as the saturated gluon system in a proton, can be connected to the concept of a saturon. This concept, introduced by Dvali et al., refers to highly occupied field configurations that reach maximal entropy while respecting unitarity. In Sec. II, we offer a concise review of QCD’s nonlinear small-$x$ evolution equations and the phenomenon of gluon saturation within the proton. To assess the saturon status of the proton, we adopt the thermodynamic entropy definition for gluonic systems as originally proposed by Kutak in Ref. [14]. Our work focuses on applying this entropy measure—along with the gluon occupation number derived from the Balitsky-Kovchegov (BK) equation—to the specific case of proton and nuclear targets.In Sec. III, we present our computed observables and examine the relationship between high-gluon-occupancy protons and saturon behavior. Finally, we provide a summary and outlook. This study introduces a novel perspective on saturation in the proton system, suggesting that, near the saturation threshold, universal behavior can emerge independently of detailed microdynamics or initial conditions. Such a proton state represents a saturated condition comprised of highly occupied soft quanta, sufficient to form a cross section constrained by unitarity [3, 4, 12].
@@ II.1 Gluon saturation and small-$x$ evolution
The unbounded growth of high-energy gluons violates unitarity. In practice, as the number of gluons increasingly populates the phase space of the proton, spatial overlap among them becomes inevitable. When the overlapping gluons come sufficiently close, gluon recombination occurs, leading to saturation (See FIG. 1). This nonlinear effect is not accounted for within the framework of the BFKL equation. In fact, the earliest nonlinear corrections to the DGLAP equation were introduced in Refs. [43, 44] and are known as the GLR-MQ equation. Subsequently, the Jalilian-Marian-Iancu-McLerran-Weigert-Leonidov-Kovner (JIMWLK) equation [45, 46, 47, 48, 49] was developed to describe the evolution of the full weight function of the color glass condensate. The BK equation [50, 51, 52] can be derived as a mean-field approximation of the JIMWLK hierarchy, formulated based on the unitarity of the dipole scattering amplitude.
@@ II.2 Balitsky-Kovchegov equation
In this work, we employ the BK equation to characterize the saturation behavior of the proton and compare it with the criteria for a saturon system. In this work, we employ the BK equation to characterize the saturation behavior of the proton. Let $\mathcal{N}(k,Y)$ denote the dipole scattering amplitude in momentum space, which depends on the transverse momentum $k$ and the evolution rapidity $Y=\ln(1/x)$. Equation (1) represents the momentum-space formulation of the BK equation, which is related to the coordinate-space version in Refs. [50, 51] via a Fourier transform (see, e.g., Ref. [53]). It is expressed as: [50, 51]
@@
While these two forms are mathematically equivalent representations of the same physical evolution, for our analysis we explicitly treat the dipole amplitude $\mathcal{N}(k,Y)$ as a function of the evolution rapidity $Y$ and the gluon transverse momentum $k$. Both formulations of the BK equation admit solutions in momentum space and can serve as inputs for calculating scattering amplitudes in diffractive processes as well as proton structure functions. Once the dipole amplitude is obtained from solving the BK equation, one can proceed to compute the unintegrated gluon distribution (UGD) and the gluon number density. Therefore, the solution to the BK equation captures the gluon evolution dynamics in the small-$x$ regime. For the nonlinear integro-differential form of Eq. (1), a Runge-Kutta scheme incorporating integral terms can be employed. In this work, we use the BKsolver algorithm as outlined in Ref. [53], which approximates the integral term using Chebyshev polynomials. This method proves highly effective and allows for both fixed and running coupling scenarios. For the differential form of the BK equation, it is also possible to obtain approximate analytical solutions. By transforming the equation into the form of the Fisher-Kolmogorov-Petrovsky-Piskunov (FKPP) equation [54, 55, 56, 57, 58, 59, 60], one can apply techniques developed for nonlinear partial differential equations to perform the analysis.
@@
For the case of running coupling (rc), the details of the treatment using the homogeneous balance method can similarly be found in Ref. [64]. In this work, we adopt the saturation scale prescription (SSP). We note that while a rigorous derivation of the running coupling in small-$x$ evolution requires a loop computation to resum vacuum polarization bubbles as established by Balitsky [65], the SSP serves here as a phenomenological ansatz whereby the running coupling scale is identified with the saturation scale $Q_{s}$. This choice reflects the characteristic momentum of gluons in the saturated regime and offers improved performance in describing the small-$x$ proton structure function $F_{2}$. The scale-dependent running coupling constant is expressed as [64]
@@
where $b=(11N_{c}-2N_{f})/12N_{c}$. Note that this coefficient differs from the standard QCD $\beta$-function coefficient by a factor of $N_{c}/\pi$, arising from the definition of the reduced coupling $\bar{\alpha}_{s}=(N_{c}/\pi)\alpha_{s}$.
@@
with the free parameter $c$. The transition from the exponential growth $e^{\lambda Y}$ in the fixed coupling case to the slower $e^{\sqrt{Y}}$ behavior in the running coupling scenario has a clear physical origin. As the saturation scale $Q_{s}(Y)$ increases with rapidity, the effective coupling $\bar{\alpha}_{s}(Q_{s}^{2})$ decreases. This dynamical reduction in the coupling strength suppresses the gluon emission probability, thereby slowing down the evolution speed of the saturation wavefront compared to the fixed coupling case. This is consistent with the reasoning behind Eq. (5), where, upon assuming a specific scale for the coupling constant, the analytical solution to the BK equation in the running coupling scenario can also be obtained using the homogeneous balance method. The explicit form is given as [64]
@@
A detailed discussion of the numerical solutions to the BK equation is beyond the scope of this work; here we briefly outline the method for extracting the saturation scale $Q_{s}^{2}$ from such solutions. While the standard definition of the saturation scale is typically formulated in coordinate space (as the inverse radius $r=1/Q_{s}$ where the amplitude reaches a fixed value), it is operationally natural in our momentum-space analysis to extract $Q_{s}$ directly from the momentum-dependent amplitude. We verify that for appropriate threshold choices (e.g., $\kappa\sim 0.5$), this momentum-space extraction reproduces the correct universal asymptotic behavior at large rapidity, consistent with theoretical expectations derived in coordinate space [55]. Notably, for $\kappa=1$, the numerical solutions obtained using BKsolver [53] reproduce the expected asymptotic behavior well in both fixed and running coupling scenarios.
@@ II.3 Saturons, occupation number and entropy
$N_{g}(x,\boldsymbol{k}_{\perp},\boldsymbol{x}_{\perp})\equiv\frac{(2\pi)^{3}}{2(N_{c}^{2}-1)}\mathrm{\penalty 10000\ }x\frac{\mathrm{d}N_{\mathrm{gluon}}}{\mathrm{d}x\mathrm{\penalty 10000\ d}^{2}\boldsymbol{k}_{\perp}\mathrm{d}^{2}\boldsymbol{x}_{\perp}},$ (10)
where $\boldsymbol{k}_{\perp}$ and $\boldsymbol{x}_{\perp}$ are transverse momentum and impact parameter of gluon, respectively. Eq. (10) defines the microscopic gluon occupation number in phase space. The standard collinear gluon distribution $xG(x,Q^{2})$ is obtained by integrating this density:
$xG(x,Q^{2})=\int^{Q^{2}}\frac{d^{2}k_{\perp}}{(2\pi)^{2}}\int d^{2}b_{\perp}\,2(N_{c}^{2}-1)N_{g}(x,k_{\perp},b_{\perp}).$ (11)
Assuming a homogeneous transverse density profile within a proton of radius $R$, the integration over impact parameter yields a geometric factor of $\pi R^{2}$. We can thus invert this relation to define an average gluon occupation number:
$N_{g}(x,Q^{2})\simeq\frac{xG(x,Q^{2})}{\pi R^{2}Q^{2}}.$ (12)
Regarding the unintegrated gluon distribution (UGD) $\mathcal{F}(x,k^{2})$, we adopt a form inspired by Ref. [70]:
$\mathcal{F}(x,k^{2})=\frac{N_{c}R^{2}}{\alpha_{s}\pi}\left(1-k^{2}\frac{d}{dk^{2}}\right)^{2}k^{2}\mathcal{N}(k,Y).$ (13)
We note that while Ref. [70] utilizes an auxiliary function without the explicit area factor, we include the term $\pi R^{2}$ here to ensure dimensional consistency with the definition of the average occupation number in Eq. (12).
@@
For simplicity, the lowest entropy state $S_{0}$ can be set to zero. All integration constants can be absorbed into quantities related to the gluon occupation number, which may be determined from experimental data or global QCD fits. It is important to emphasize that, in this analysis, the gluon occupation number (or gluon density) is obtained from the solution of the BK equation. Ultimately, following the thermodynamic definition proposed by Kutak in Ref. [14], the entropy of the system takes the form:
@@ II.4 From proton to nucleus: Glauber-Gribov-Mueller Scheme
where $Q_{s,p}=Q_{s}$ in Eq. (5) represents the nucleon saturation scale, $R_{A}$ is the nuclei radius with mass number $A$ and $r_{0}=1.2$ fm. We note that the scaling $Q_{s,A}^{2}\sim A^{1/3}$ assumes a simplified cylindrical geometry corresponding to the maximum optical depth in central collisions. As pointed out in recent phenomenological studies [90, 91], a more realistic treatment involving the Woods-Saxon nuclear density profile and averaging over impact parameters (particularly for minimum-bias events) leads to a significantly more modest growth of the saturation scale. Therefore, the estimate in Eq. (20) should be regarded as an upper bound on the nuclear enhancement. However, even with a softer scaling, the saturation scale of a heavy nucleus remains substantially larger than that of a proton, preserving the qualitative ordering discussed in this work.
@@ III.1 Proton case
Through the mean-field approximation of the nonlinear evolution equation, namely the momentum-space BK equation, both its analytical and numerical solutions for the dipole amplitude can be related to the gluon distribution via the definition of the UGD. This connection allows the determination of the gluon occupation number and the corresponding thermodynamic entropy, as given in Eqs (12-18).
@@
The parameter choices for all results shown in the figure are specified in the main text and are taken from the corresponding references. The four subplots, arranged from left to right and top to bottom, represent respectively: the analytical solution of the BK equation with fixed coupling, the analytical solution with running coupling, the numerical solution with fixed coupling, and the numerical solution with running coupling. Each panel displays the resulting gluon occupation number and the associated entropy. It is observed that in all cases, even down to $x=10^{-7}$—well below the kinematic reach of current DIS experiments—neither the entropy nor the gluon occupation number approaches the inverse coupling constant, $1/\alpha_{s}$. We note that while the figures include trajectories extending to $x\sim 10^{-1}$ to visualize the initialization and continuity of the evolution, our physical conclusions regarding the saturon bound are strictly based on the behavior in the deep small-$x$ region ($x<10^{-3}$) where the BK formalism is theoretically valid. One important reason for this is the presence of approximations in the construction of the entropy, and the omission of higher-order corrections in the evolution equations.
Regarding the sensitivity to initial conditions, it is well known that the absolute value of the saturation scale $Q_{s}$ depends on the parameters of the initial distribution at $Y=0$. However, a key feature of the BK evolution is the emergence of traveling wave solutions, where the asymptotic velocity of the wavefront is determined by the evolution kernel rather than the initial inputs. This implies that while varying the initial conditions might horizontally shift the saturation scale (and consequently the entropy curves), the rate of growth with rapidity is a robust prediction of the theory. Therefore, our conclusion that the growth rate is insufficient to reach the saturon limit within the accessible kinematic window remains valid across reasonable variations of the initial parameters.
@@ III.2 Nuclear case
As seen in the figure, compared with the proton result, under the $x$-dependent growth of the entropy, the thermodynamic entropy of a highly occupied lead nucleus reaches the saturation criterion ($1/\alpha_{s}$) at around $x\sim 10^{-6}$. This indicates that the nucleus indeed enhances the entropy, mainly through the increase of the transverse radius after the high-energy boost (set by the nuclear radius) and the amplification of the nuclear saturation scale. However, this comparison has limitations, such as the choice of the dipole-amplitude form and uncertainties from nuclear geometric parameters. In this section, we present results using fixed coupling. While incorporating running coupling effects would suppress the evolution speed (as seen in the proton case), it would not alter the qualitative ordering between the proton and the nucleus. The primary advantage of the nuclear target arises from the initial enhancement factor $Q_{s,A}^{2}\simeq A^{1/3}Q_{s,p}^{2}$. For a lead nucleus ($A=208$), this provides a large “geometric head start.” Even if the subsequent evolution were slowed down by running coupling corrections, the nuclear entropy curve would remain vertically shifted significantly above the proton curve throughout the relevant kinematic window. Thus, the conclusion that nuclei approach the saturon bound more readily than protons is robust against higher-order corrections.